RADEMACHER’S THEOREM ON CONFIGURATION SPACES AND APPLICATIONS

MICHAEL RÖCKNER AND ALEXANDER SCHIED

Abstract. We consider an L^{2}-Wasserstein type distance ρ on the configuration space Γ_{X} over a Riemannian manifold X, and we prove that ρ-Lipschitz functions are contained in a Dirichlet space associated with a measure on Γ_{X} satisfying some general assumptions. These assumptions are in particular fulfilled by a large class of tempered grandcanonical Gibbs measures with respect to a superstable lower regular pair potential. As an application we prove a criterion in terms of ρ for a set to be exceptional. This result immediately implies, for instance, a quasi-sure version of the spatial ergodic theorem. We also show that ρ is optimal in the sense that it is the intrinsic metric of our Dirichlet form.